# Detecting the tidal heating with the generic extreme-mass-ratio inspirals

### v2

@@ III.2 Non-spinning massive objects
In the subsection we present similar results by employing EMRI waveforms from non-spinning ECOs. The EMRI trajectories modified by tidal heating are computed utilizing the analytic fluxes in the Schwarzschild spacetime Munna et al. (2020, 2023).
Since the azimuthal, radial and polar dephasings exhibit the similar behaviors, except for the difference in magnitudes, here we only show the azimuthal dephasing as a function of observation time in Fig. 4. The horizontal black and dashed line denote the threshold value that can be distinguished by LISA. For the EMRI waveforms from the ECO with $|\mathcal{R}|^{2}=10^{-3}$ and the initial eccentricity $e_{0}=0.5$, the tidal heating effect can be detected with the observation of LISA of four months.
Figure 4: The azimuthal dephasing $\delta\Phi_{\phi}$ as a function of the observation time is plotted. The parameters are taken as $p_{0}=12,e_{0}=0.5,x_{I}=0.5,|\mathcal{R}|^{2}\in(10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1})$.
Figure 5: Mismatch as a function of observation time for EMRI waveforms from ECO and Schwarzschild BH is plotted, where the other parameters are set as $p_{0}=12,x_{I}=0.5$. The left panel shows the effects from various values of the reflectivity parameter with $e_{0}=0.7$, and the right panel focuses the effects from various values of the orbital eccentricity with$|\mathcal{R}|^{2}=10^{-3}$.
Then we also analyze the mismatch of the two EMRI waveforms from the Schwarzschild BH and the non-spinning ECO cases in Fig. 5. The black dashed line corresponds to the threshold value of mismatch i.e., $\mathcal{M}_{\rm min}=0.001$. From the left panels of Fig. 3 and Fig. 5, one can see that with the same sources parameters, the mismatch in the non-spinning case is slightly smaller than that in the spinning case. For example, for the ECO with a reflectivity of $|\mathcal{R}|^{2}=5\times 10^{-4}$ and an initial eccentricity of $e_{0}=0.5$, it requires approximately three months to attain the threshold mismatch value in the spinning scenario. In contrast, the corresponding period extends to about five months in the non-spinning case. This implies that the spin can enhance the detectability of the tidal heating effect. From the right panel of Fig. 5, as a whole, the mismatch grows with the initial eccentricity. For an initial eccentricity $e_{0}=0.5$, LISA can not discern the tidal heating effect with one year observation for the non-spinning primaries, however, the observation of four months can be distinguished for the spinning primaries in the right panel of Fig. 3.
@@
$p_{0}$ $e_{0}$ $\Delta M/M$ $\Delta m/m$ $\Delta a$ $\Delta e_{0}$ $\Delta p_{0}$ $\Delta(|\mathcal{R}|^{2})$ $\Delta D_{L}/D_{L}$ $\Delta\Omega_{S}$ $\Delta\Omega_{K}$
$12$ $0.4$ $2.93\times 10^{-4}$ $7.04\times 10^{-4}$ $5.21\times 10^{-4}$ $7.68\times 10^{-5}$ $9.54\times 10^{-4}$ $2.25\times 10^{-3}$ $7.58\times 10^{-2}$ $6.09\times 10^{-4}$ $7.86$
$12$ $0.8$ $9.91\times 10^{-5}$ $8.28\times 10^{-5}$ $1.59\times 10^{-4}$ $6.45\times 10^{-5}$ $8.71\times 10^{-4}$ $5.42\times 10^{-4}$ $4.65\times 10^{-2}$ $1.44\times 10^{-3}$ $5.62$
Table 2: Measurement errors of the full EMRI source parameters are listed, the other parameters are set as $\Phi_{r}^{0}=1.0$, $\Phi_{\theta}^{0}=1.0$, $\Phi_{\phi}^{0}=1.0$ and $|\mathcal{R}|^{2}=0.001$.
@@ IV Conclusion
A key and interesting goal of the EMRI detections is to probe the nature of dark compact objects Cardoso and Pani (2019). Recent studies have developed proposals to distinguish massive horizonless objects based on considerations of tidal heating Datta and Bose (2019); Datta et al. (2020), tidal deformability Pani and Maselli (2019); De Luca et al. (2023); Zi and Li (2023) and area quantization Agullo et al. (2021); Datta and Phukon (2021). In this paper we analyzed the detectability of waveforms from the generic EMRIs around the ECOs, then calculated the constraint on the reflectivity parameter with the FIM method.
Utilizing precise analytic flux formulas derived for Kerr Sago and Fujita (2015) and Schwarzschild BHs Munna et al. (2020, 2023), we adapted these formulas in the vicinity of the ECO surface by incorporating a reflectivity parameter, employing the methodology outlined in Ref. Datta et al. (2020). Injecting the modified fluxes into the few model, we computed the EMRI waveforms of the ECOs. Then we analyzed the differences of EMRI waveforms from the ECO and the Kerr or Schwarzschild BH spacetime by computing the dephasing and mismatch. Our results indicate that one year observation of EMRI by LISA can distinguish the tidal heating corrected signals with a reflectivity as smaller as $|\mathcal{R}|^{2}\sim 10^{-4}$ for the spinning ECO, however, the EMRI signal from the non-spinning ECO can constrain the reflectivity parameter at the level of $|\mathcal{R}|^{2}\sim 5\times 10^{-4}$. Finally, we assessed the capability of LISA to constrain the tidal heating effect using the FIM method. The measurement accuracy of the reflectivity of the ECOs through LISA observations can achieve a precision on the order of $\mathcal{O}(10^{-4})$. The constraint level is expected to be further enhanced with the incorporation of the orbital eccentricity.
@@ Acknowledgments
The work is in part supported by NSFC Grant No.12205104, “the Fundamental Research Funds for the Central Universities” with Grant No. 2023ZYGXZR079, the Guangzhou Science and Technology Project with Grant No. 2023A04J0651 and the startup funding of South China University of Technology. T. Z. is funded by China Postdoctoral Science Foundation Grant No. 2023M731137. The work of T. Z. is also supported by the National Natural Science Foundation of China (Grants No. 12347140).

### v6

@@ III.2 Non-spinning massive objects
In the subsection we present similar results by employing EMRI waveforms from non-spinning ECOs. The EMRI trajectories modified by the ECOs are computed utilizing the analytic fluxes in the Schwarzschild spacetime Munna et al. (2020, 2023).
Since the azimuthal, radial and polar dephasings exhibit the similar behaviors, except for the difference in magnitudes, here we only show the azimuthal dephasing as a function of observation time in Fig. 5. The horizontal black dashed line denotes the threshold value that can be distinguished by LISA. For the EMRI waveforms from the ECO with $|\mathcal{R}|^{2}=10^{-3}$ and the initial eccentricity $e_{0}=0.5$, the tidal heating effect can be discerned with the observation of LISA of four months.
Figure 5: The azimuthal dephasing $\delta\Phi_{\phi}$ as a function of the observation time is plotted. The parameters are taken as $p_{0}=12,e_{0}=0.5,\theta_{I}=\pi/3,|\mathcal{R}|^{2}\in(10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1})$.
Figure 6: Mismatch as a function of observation time for EMRI waveforms from ECO and Schwarzschild BH is plotted, where the other parameters are set as $p_{0}=12,\theta_{I}=\pi/3$. The left panel shows the effects from various values of the reflectivity parameter with $e_{0}=0.7$, and the right panel focuses the effects from various values of the orbital eccentricity with$|\mathcal{R}|^{2}=10^{-3}$.
Then we also analyze the mismatch of the two EMRI waveforms from the Schwarzschild BH and the non-spinning ECO cases in Fig. 6. From the left panels of Fig. 3 and Fig. 6, one can see that with the same sources parameters, the mismatch in the non-spinning case is slightly smaller than that in the spinning case. From the right panel of Fig. 6, as a whole, the mismatch grows with the initial eccentricity. For an initial eccentricity $e_{0}=0.5$, LISA can not discern the difference of the tidal heating effect with one year observation for the non-spinning primaries, however, the observation of four months can be distinguished for the spinning primaries in the right panel of Fig. 3.
@@
$p_{0}$ $e_{0}$ $\Delta\ln M$ $\Delta\ln m$ $\Delta e_{0}$ $\Delta p_{0}$ $\Delta|\mathcal{R}|^{2}$ $\Delta D_{L}/D_{L}$ $\Delta\Omega_{S}$ $\Delta\Omega_{K}$
$12$ $0.4$ $1.19\times 10^{-4}$ $2.44\times 10^{-3}$ $1.61\times 10^{-3}$ $9.18\times 10^{-4}$ $1.03\times 10^{-3}$ $8.67\times 10^{-2}$ $7.59\times 10^{-3}$ $7.86$
$12$ $0.8$ $7.51\times 10^{-5}$ $8.28\times 10^{-4}$ $1.45\times 10^{-3}$ $8.71\times 10^{-4}$ $5.42\times 10^{-4}$ $4.65\times 10^{-2}$ $5.54\times 10^{-3}$ $5.62$
Table 2: Measurement errors of the full EMRI source parameters are listed, the other parameters are set as $\Phi_{r}^{0}=1.0$, $\Phi_{\phi}^{0}=1.0$ and $|\mathcal{R}|^{2}=0.001$.
@@
Figure 7: The corner plot for the probability distribution for the mass of secondary body, mass of primary body, initial orbital eccentricity, initial semi-latus rectum and reflectivity $(M,m,e_{0}=0.4,p_{0}=12,|\mathcal{R}|^{2})$, which is inferred from one-year observation of EMRI. Vertical lines show the $1\sigma$ interval for each source parameter. The contours correspond to $68\%$, $95\%$ and $99\%$ probability confidence intervals
In the end, we also argue the correlations of constraint reflectivity among the other intrinsic parameters in the non-spinning primary case. We display the probability distribution for the mass of secondary body, mass of primary body, initial orbital eccentricity, initial semi-latus rectum and reflectivity $(M,m,e_{0},p_{0},|\mathcal{R}|^{2})$ in Fig. 7. One can find that the bound on the reflectivity is closely associated with the parameters $(e_{0},p_{0})$, and the correlation between the reflectivity $|\mathcal{R}|^{2}$ and celestial bodies $(m,M)$ is relative weaker.
@@ IV Conclusion
A key and interesting goal of the EMRI detections is to probe the nature of dark compact objects Cardoso and Pani (2019). Recent studies have developed proposals to distinguish massive horizonless objects based on considerations of tidal heating Datta and Bose (2019); Datta et al. (2020), tidal deformability Pani and Maselli (2019); De Luca et al. (2023); Zi and Li (2023) and area quantization Agullo et al. (2021); Datta and Phukon (2021). In this paper we analyzed the difference of waveforms from the generic EMRIs around the ECOs, then calculated the constraint on the reflectivity parameter with the FIM method.
Utilizing precise analytic flux formulas derived for Kerr Sago and Fujita (2015) and Schwarzschild BHs Munna et al. (2020, 2023), we adapted these formulas in the vicinity of the ECO surface by incorporating a reflectivity parameter, employing the methodology outlined in Ref. Datta et al. (2020). Injecting the modified fluxes into the FEW model, we computed the EMRI waveforms of the ECOs. Then we analyzed the differences of EMRI waveforms from the ECO and the Kerr or Schwarzschild BH spacetime by computing the dephasing and mismatch. Our results indicate that one year observation of EMRI by LISA can distinguish the tidal heating corrected signals with a reflectivity as smaller as $|\mathcal{R}|^{2}\sim 10^{-6}$ for the spinning ECO, however, the EMRI signal from the non-spinning ECO can constrain the reflectivity parameter at the level of $|\mathcal{R}|^{2}\sim 5\times 10^{-4}$. Finally, we assessed the capability of LISA to constrain the tidal heating effect using the FIM method. The measurement accuracy of the reflectivity of the ECOs through LISA observations can achieve a precision on the order of $\mathcal{O}(10^{-6})$. The constraint level is expected to be further enhanced with the incorporation of the orbital eccentricity. According to Figs. 4 and 7, Note that the reflectivity $|\mathcal{R}|^{2}$ and the other parameters $(M,a,e_{0},p_{0},x_{I})$ are correlated for the rotating ECO case, the correlations between reflectivity $|\mathcal{R}|^{2}$ and the other parameters $(M,e_{0},p_{0})$ are positive for the non-spinning ECO case. In summary, the reflectivity of the spinning ECO could be more rigorously constrained using eccentric EMRI signals compared to the non-spinning ECO case.
@@ Acknowledgments
The work is in part supported by NSFC Grant No.12205104, “the Fundamental Research Funds for the Central Universities” with Grant No. 2023ZYGXZR079, the Guangzhou Science and Technology Project with Grant No. 2023A04J0651 and the startup funding of South China University of Technology. T. Zi is funded by the China Postdoctoral Science Foundation with Grant No. 2023M731137 and the National Natural Science Foundation of China with Grants No. 12347140. T.Zi thanks Chao Zhang for very helpful discussions on the corner plots.